OpenStudy (anonymous):

Solve using the multiplication principle. -3/7x=6/35

OpenStudy (mertsj):

Multiply both sides by the reciprocal of - 3/7

OpenStudy (anonymous):

so you multiply both sides by -3/7?

OpenStudy (mertsj):

What is the reciprocal of -3/7?

OpenStudy (anonymous):

not sure :/

OpenStudy (mertsj):

Could you do this problem? \[\frac{3}{4}\div\frac{-9}{16}\]

OpenStudy (anonymous):

is it -1/2?

OpenStudy (mertsj):

My point is, you have to multiply by the reciprocal. What is the reciprocal of -9/16?

OpenStudy (mertsj):

\[\frac{3}{4}\times\frac{-16}{9}\]